OpenStudy (anonymous):

8m+n=34 m-3n=23 Simplify using substitution method

OpenStudy (anonymous):

i think the first one is m+n =4.25

OpenStudy (anonymous):

solve the second equation for m and put that in for m of the first equation. Distribute and solve and you'll get n. Then plug your n value into either equation and solve for m.

OpenStudy (precal):

You can use m=3n+23

OpenStudy (precal):

Sub into 8m+n=34 8(3n+23)+n=34 solve for n